# Active Advantage-Aligned Online Reinforcement Learning with Offline Data

### v3

@@ Abstract
Online reinforcement learning (RL) enhances policies through direct interactions with the environment, but faces challenges related to sample efficiency. In contrast, offline RL leverages extensive pre-collected data to learn policies, but often produces suboptimal results due to limited data coverage. Recent efforts integrate offline and online RL in order to harness the advantages of both approaches. However, effectively combining online and offline RL remains challenging due to issues that include catastrophic forgetting, lack of robustness to data quality and limited sample efficiency in data utilization. In an effort to address these challenges, we introduce $A^{3}$RL, which incorporates a novel confidence-aware Active Advantage-Aligned ($A^{3}$) sampling strategy that dynamically prioritizes data aligned with the policy’s evolving needs from both online and offline sources, optimizing policy improvement. Moreover, we provide theoretical insights into the effectiveness of our active sampling strategy and conduct diverse empirical experiments and ablation studies, demonstrating that our method outperforms competing online RL techniques that leverage offline data.
Introduction
@@
Reinforcement learning (RL) has achieved notable success in many domains, such as robotics (Kober, Oztop, and Peters 2011; Kober, Bagnell, and Peters 2013), game play (Mnih et al. 2013; Silver et al. 2017), drug discovery (Liu et al. 2023a, 2024b), and reasoning with Large Language Models (LLMs) (Havrilla et al. 2024). Online RL algorithms such as Q-learning (Watkins 1989), SARSA (Rummery and Niranjan 1994), and PPO (Schulman et al. 2017) learn and make decisions in an online, sequential manner, whereby an agent interacts with an environment and learns from its experience. However, due to the need for exploration that is fundamental to RL, online RL tends to be highly sample inefficient in high-dimensional or sparse reward environments. A complementary approach to improve the sample efficiency is imitation learning (IL)  (Ross and Bagnell 2010, 2014), where an agent learns a policy by leveraging expert demonstrations (Cheng, Kolobov, and Agarwal 2020; Liu et al. 2023b, c).
@@
However, in many cases, we do not have access to a live expert to query, but often have access to an abundance of logged data collected from experts. One approach to make use of this data is through offline reinforcement learning. Offline RL (Levine et al. 2020; Prudencio, Maximo, and Colombini 2023) learns a policy solely from such a fixed dataset of pre-collected experiences, without the need to directly interact with the environment. Despite its advantages, offline RL often results in a suboptimal policy due to dataset limitations. This has motivated recent work that combines offline and online RL, whereby learning begins from a logged dataset before transitioning to online interactions for further improvement. While beneficial, contemporary offline-to-online RL methods suffer from catastrophic forgetting, where previously learned knowledge is overwritten during online fine-tuning, leading to significant performance degradation (Luo et al. 2023; Zheng et al. 2023).
@@
More recently, methods that integrate online RL with offline datasets utilize off-policy techniques to incorporate offline data while learning online (Ball et al. 2023; Song et al. 2022), mitigating catastrophic performance drops. These techniques do not require any preliminary offline RL training or incorporate specific imitation clauses that prioritize pre-existing offline data. Notably, RLPD (Ball et al. 2023) exhibits strong empirical performance, however it employs a uniform random sampling strategy for both offline and online learning, ignoring that different transitions contribute differently to the various stage of policy improvement. Furthermore, this uniform sampling strategy may result in data inefficiencies (e.g., sampling useless data while missing valuable data) and also make policy improvement highly sensitive to data quality.
Figure 1: Method Overview
Our contributions.
@@
In this work, we introduce Active Advantage-Aligned Reinforcement Learning ($A^{3}$RL), a novel method that operates in the realm of online RL with an offline dataset, as illustrated in Fig. 1. Our approach dynamically prioritizes data (transitions) that have the highest potential to maximize policy improvement, aligning with the evolving quality and learning needs of the policy. More specifically, $A^{3}$RL considers not only the relevance of the data in facilitating the current policy’s online exploration and exploitation but also its estimated contribution to policy improvement via confidence-aware advantage-based prioritization. $A^{3}$RL demonstrates robustness to data quality in a black-box manner and maintains resilience under varying environmental conditions. Notably, it also effectively accelerates policy improvement, even in a purely online environment.
In summary, our contributions are:
@@
  • We propose $A^{3}$RL, a novel algorithm for online RL with offline data. This algorithm surpasses current state-of-the-art (SOTA) methods by integrating a priority-based active sampling strategy based on the value of confidence-aware advantage function and coverage by offline dataset.
@@
  • In contrast to RLPD and other related works (Lee et al. 2022; Schaul et al. 2015), which lack theoretical support, this study provides theoretical insights of our confidence-aware active advantage-aligned sampling strategy, demonstrating superiority and its minimum improvement gap over random sampling.
@@
  • Through extensive empirical evaluations in various environments, we demonstrate that $A^{3}$RL achieves consistent and significant improvements over prior SOTA models.
@@
  • Given the black-box nature of offline datasets, we conduct comprehensive ablation studies across a range of dataset qualities and environmental settings, including purely online scenarios, to evaluate the robustness of $A^{3}$RL. These studies consistently confirm its stable performance across diverse conditions, regardless of environmental factors or data quality.
Related Work
Online RL with offline datasets
@@
Several methods exist that incorporate offline datasets in online RL to enhance sample efficiency. Many rely on high-quality expert demonstrations (Ijspeert, Nakanishi, and Schaal 2002; Kim et al. 2013; Nair et al. 2018; Rajeswaran et al. 2017; Vecerik et al. 2017; Zhu et al. 2019). Nair et al. (2020) introduced the Advantage Weighted Actor Critic (AWAC), which utilizes regulated policy updates to maintain the policy’s proximity to the observed data during both offline and online phases. On the other hand, Lee et al. (2022) propose an initially pessimistic approach to avoid over-optimism and bootstrap errors in the early online phase, gradually reducing the level of pessimism as more online data becomes available. Most relevant to our work is RLPD (Ball et al. 2023), which adopts a sample-efficient off-policy approach to learning that does not require pre-training. Unlike RLPD, which utilizes symmetric sampling to randomly draw from both online and offline datasets for policy improvement, $A^{3}$RL adopts a Prioritized Experience Replay (PER)-style method, whereby it selectively uses data from both datasets to enhance policy performance.
Prioritized experience replay.
@@
Experience replay (Lin 1992) enhances data efficiency in online RL by reusing past experiences. Priority Experience Replay (PER) (Schaul et al. 2015) introduces prioritization based on temporal difference (TD) error to ensure that impactful experiences are used more frequently, and has proven effective in a variety of settings (Hessel et al. 2018; Jaderberg et al. 2016; Nair et al. 2018; Oh et al. 2021; Saglam et al. 2022; Tian et al. 2023; Wang et al. 2016; Yue et al. 2023). Alternative prioritization strategies have been explored, such as prioritizing transitions based on expected return (Isele and Cosgun 2018) or adjusting sample importance based on recency (Fedus et al. 2020). Existing research predominantly focuses on either purely online or offline applications of PER. Our research distinctively integrates the advantages of both online and offline data in an innovative way. Eysenbach et al. (2020) apply a density ratio to the reward instead of weighting the samples. The most relevant studies to ours include Sinha et al. (2022) that uses the density ratio between off-policy and near-on-policy state-action distributions as an importance weight for policy evaluation, and Lee et al. (2022) that employs density ratios to select relevant samples from offline datasets. Our method differs by not only using the density ratio to assess the “on-policyness” of the data but also by considering the confidence-aware advantage value to determine how much the data can contribute to enhancing policy improvement.
Active learning in RL.
@@
Active learning has been explored in RL for data-efficient exploration (Epshteyn, Vogel, and DeJong 2008; Fang, Li, and Cohn 2017; Krueger et al. 2020; Liu et al. 2022, 2023b, 2023c; Lopes, Melo, and Montesano 2009). Unlike previous approaches that focus on oracle selection (Liu et al. 2023b, c), state exploration (Epshteyn, Vogel, and DeJong 2008; Liu et al. 2023c) or reward estimation (Lopes, Melo, and Montesano 2009), $A^{3}$RL introduces an active transition sampling mechanism tailored to online RL with offline data, prioritizing transitions that maximize policy improvement. We defer more details of related work to Appendix.
Preliminaries and Problem Statement
@@
We consider a discounted Markov decision process (MDP) environment (Bellman 1957) characterized by a tuple $M=(\mathcal{S},\mathcal{A},\mathcal{P},R,\gamma,d_{0})$, where $\mathcal{S}$ represents a potentially infinite state space, $\mathcal{A}$ is the action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the unknown transition kernel, $R:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the reward function, $\gamma\in(0,1)$ is the discount factor and $d_{0}\left({s}\right)$ is the initial state distribution. The learner’s objective is to solve for the policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ that maximizes the expected sum of discounted future rewards $\mathbb{E}_{\pi}\!\left[\sum_{t=1}^{\infty}\gamma^{t}r\left({s_{t},a_{t}}\right)\right]$, where the expectation is taken over the trajectory sampled from $\pi$.
Maximum entropy RL.
@@
In this work, we adopt off-policy soft actor-critic (SAC) (Haarnoja et al. 2018) RL to train an agent with samples generated by any behavior policy. We use a general maximum entropy objective (Ball et al. 2023; Haarnoja et al. 2018; Ziebart 2010) as follows:
$\max_{\pi}\mathbb{E}_{s\sim\rho^{\pi},{a\sim\pi}}\!\left[\sum_{t=0}^{\infty}\gamma^{t}\left({r_{t}+\alpha\mathcal{H}\left({\pi\left({a|s}\right)}\right)}\right)\right],$ (1)
@@
where $\alpha$ is a temperature parameter. This involves optimizing reward while encouraging exploration, making the learned policy more robust.
$Q$-value and advantage function.
@@
The $Q$-value function measures the expected return of executing action $a$ in state $s$ under policy $\pi$: $Q^{\pi}\left({s,a}\right)=\mathcal{B}^{\pi}Q^{\pi}\left({s,a}\right)$, where $\mathcal{B}^{\pi}$ is the Bellman operator: $\mathcal{B}^{\pi}Q\left({s,a}\right):=r\left({s,a}\right)+\gamma\mathbb{E}_{s^{\prime}\sim P\left({\cdot|s,a}\right)}\!\left[V^{\pi}\left({s^{\prime}}\right)\right]$. The soft state value function is defined as: $V^{\pi}\left({s}\right):=\mathbb{E}_{a\sim\pi\left({\cdot|s}\right)}\!\left[Q^{\pi}\left({s,a}\right)-{\log\pi\left({a|s}\right)}\right]$. For a generator policy $\pi$, the advantage function (Sutton et al. 1999) quantifies the relative benefit of selecting $a$ over the policy’s default behavior:
@@
Specifically, SAC learns a soft Q-Function, denoted as $Q_{\theta}\left({s,a}\right)$ which parameterized by $\theta$, and a stochastic policy $\pi_{\phi}$ parameterized by $\phi$. The SAC method involves alternating between updates for the critic and the actor by minimizing their respective objectives (Lee et al. 2022) as follows
@@
$-\gamma{\mathbb{E}_{a_{t+1}\sim\pi_{\phi}}\!\left[Q_{\overline{\theta}}\left({s_{t+1},a_{t+1}}\right)-\alpha\log\pi_{\phi}\left({a_{t+1}|s_{t+1}}\right)\right]})^{2}],$
@@
Here, $\mathcal{R}$ is an experience replay buffer of either on-policy experience (Sutton et al. 1999) or through off-policy experience (Munos et al. 2016; Precup 2000), and $\overline{\theta}$ denotes the delayed parameters.
@@
Prioritized experience replay. PER (Schaul et al. 2015) serves as the basis of our sampling techniques, providing a framework for prioritizing experience replay based on transition importance. Instead of sampling uniformly from the replay buffer $\mathcal{R}$, PER assigns higher probability to more informative transitions, leading to improved sample efficiency (Hessel et al. 2018). Each transition $\mathcal{R}_{i}=(s_{i},a_{i},r_{i},s_{i+1})$ is assigned a priority $\sigma_{i}$, typically based on the TD-error: $\delta=r+\gamma V\left({s_{t+1}}\right)-V\left({s_{t}}\right)$ (Brittain et al. 2019; Hessel et al. 2018; Oh et al. 2021; Schaul et al. 2015; Van Hasselt, Hessel, and Aslanides 2019). Subsequently, the sampling approach of PER involves establishing an index set $\mathcal{I}$ within the range of $\left[|\mathcal{R}|\right]$ based on the probabilities $p_{i}$ assigned by the priority set as follows: $p_{i}=\frac{\sigma_{i}^{\mathcal{\zeta}}}{\sum_{k\in\left[|\mathcal{R}|\right]}\sigma_{k}^{\mathcal{\zeta}}}$, with a hyper-parameter $\mathcal{\zeta}>0$. To correct for sampling bias, PER applies importance sampling weights:
$u_{i}\propto\bigl{(}{1}/(|\mathcal{R}|\cdot p_{i})\bigr{)}^{\beta},$ (3)
@@
where $\beta$ anneals from $\beta_{0}\in(0,1)$ to $1$ during training to counteract bias in the learning updates, and the importance sampling weights are normalized to have maximum weight of $1$ for stability. While standard PER prioritizes TD-error, our method extends this framework to prioritize transitions based on onlineness and contribution to policy improvement.
@@
Online RL with offline datasets. In this work, we study online RL with offline datasets denoted as $\mathcal{D}$ (Ball et al. 2023). These datasets consist of a set of tuples $(s,a,r,s^{\prime})$ generated from a specific MDP. A key characteristic of offline datasets is that they typically offer only partial coverage of state-action pairs. In other words, the set of states and actions in the dataset, denoted as $\left\{\left({s,a}\right)\in\mathcal{D}\right\}$, represents a limited subset of the entire state space and action space, $\mathcal{S}\times\mathcal{A}$. Moreover, learning on the data with incomplete coverage of state-action pairs potentially results in excessive value extrapolation during the learning process for methods using function approximation (Fujimoto, Meger, and Precup 2019). Our model, based on SAC (Haarnoja et al. 2018), incorporates several effective strategies for RL with offline data, as outlined in RLPD (Ball et al. 2023). These strategies include:
Clipped Double Q-Learning: The maximization objective of Q-learning and the estimation uncertainty from value-based function approximation often leads to value overestimation (Van Hasselt, Guez, and Silver 2016). To address this problem, Fujimoto, Hoof, and Meger (2018) introduced Clipped Double Q-Learning (CDQ) as a means of mitigation. CDQ involves taking the minimum from an ensemble of two Q-functions for computing TD-backups. The targets for updating the critics are given by the equation $y=r\left({s,a}\right)+\gamma\min_{i=1,2}Q_{\theta_{i}}\left({s^{\prime},a^{\prime}}\right)$, where $a^{\prime}\sim\pi\left({\cdot|s^{\prime}}\right)$. See Appendix B for more details.
Algorithm
Confidence-aware Active Advantage-Aligned Sampling Strategy
In this study, we theoretically derived from the performance difference lemma in §Theoretical Analysis and presented active advantage-aligned strategy, a novel sampling approach for policy improvement. Here, ‘advantage’ measures the potential impact of the transition on policy improvement, while ’aligned’ assesses how well the transition aligns with the states sampled online by the current policy. This method allows for the safe utilization of online and offline samples by harnessing relevant, near on-policy offline samples that also present the potential to enhance policy improvement. For the advantage term, to enhance robustness, we use the pessimistic CDQ $Q$ estimation, while incorporating uncertainty estimation for the value function under the current policy. Specifically, we estimate both the value function $\widehat{V}$—which directly determines the estimated advantage of $\widehat{A}$—and the associated uncertainty $\widehat{C}_{A}\left({s,a}\right)$, through Monte Carlo samples of the on-policy actions. Furthermore, we extend this approach to density ratio estimation, using an ensemble of density networks to predict the density ratio $\widehat{{w}}\left({s,a}\right)$ and associated uncertainty $\widehat{C}_{w}\left({s,a}\right)$. This approach broadens the distribution of samples used for updates, centering around on-policy examples, thereby facilitating immediate value. The active advantage-aligned priority $\sigma$} and the probability p{p} are as follows:
${p}_{i}$ $=\frac{\sigma_{i}^{\mathcal{\zeta}}}{\sum_{k\in\left[|\mathcal{R}|\right]}\sigma_{k}^{\mathcal{\zeta}}},$ (4)
${\sigma}_{i}$ $={\sigma}\left({s_{i},a_{i}}\right)=\left({\mathbb{I}^{\text{off}}\underline{{w}}\left({s_{i},a_{i}}\right)+\mathbb{I}^{\text{on}}}\right)\cdot\exp\left({\mathcal{\xi}\cdot\underline{A}\left({s_{i},a_{i}}\right)}\right),$
$\underline{{w}}$ $\left({s_{i},a_{i}}\right)=\widehat{{w}}\left({s_{i},a_{i}}\right)-\widehat{C}_{w}(s_{i},a_{i}),$ (5)
$\underline{A}$ $\left({s_{i},a_{i}}\right)=\widehat{A}\left({s_{i},a_{i}}\right)-\widehat{C}_{A}(s_{i},a_{i}),$ (6)
where $\mathbb{I}^{\text{off}}$ and $\mathbb{I}^{\text{on}}$ represents the indicator of offline and online respectively, density ratio $\underline{{w}}\left({s,a}\right)$ is the LCB (Lower Confidence Bound) (Liu et al. 2023b) of density ratio, which measures the onlineness of the transition (defined in Eq. (7)) in a conservative manner, $\underline{A}\left({s,a}\right)$ is LCB of the advantage term, which assesses the potential of the transition in improving the policy and $\mathcal{\xi}>0$ representing a temperature hyperparameter associated with the advantage term, and another $\mathcal{\zeta}>0$ for the entire priority term, per the standard PER approach. This approach considers not only the on-policyness of the data but also measures how important the data contributes to the current policy improvement. The active advantage-aligned sampling strategy aims to assign greater weight to transitions that are either not well covered by the offline dataset—indicating that the state-action pair is novel to the offline policy (i.e., the density ratio is large)—or that represent good actions for maximizing cumulative reward (i.e., the advantage / $Q$ function is large).

### v4

# Advantage-Aligned Active Online Reinforcement Learning with Offline Data

@@ Abstract
Online reinforcement learning (RL) enhances policies through direct interactions with the environment, but faces challenges related to sample efficiency. In contrast, offline RL leverages extensive pre-collected data to learn policies but often yields suboptimal results due to limited data coverage and redundancy within the dataset. Recent efforts integrate offline and online RL in order to harness the advantages of both approaches. However, effectively combining online and offline RL remains challenging due to issues that include catastrophic forgetting, lack of robustness to data quality and limited sample efficiency in data utilization. To address these challenges, we introduce $A^{3}$RL, which incorporates a novel confidence-aware Active Advantage-Aligned ($A^{3}$) sampling strategy. Our approach dynamically prioritizes samples from both online and offline data sources by aligning the sampling distribution with the current policy improvement direction, thereby enabling more efficient and effective policy optimization. Moreover, we provide theoretical insights into the effectiveness of our active sampling strategy and conduct diverse empirical experiments and ablation studies, demonstrating that our method outperforms competing online RL techniques that leverage offline data.
1 Introduction
@@
Reinforcement learning (RL) has achieved notable success in many domains, such as robotics [25, 26], game play [41, 58], drug discovery [35, 38], and reasoning with Large Language Models (LLMs) [16]. Online RL algorithms such as Q-learning [68], SARSA [54], and PPO [57] learn and make decisions in an online, sequential manner, whereby an agent interacts with an environment and learns from its experience. However, due to the need for exploration that is fundamental to RL, online RL tends to be highly sample inefficient in high-dimensional or sparse reward environments. A complementary approach to improve the sample efficiency is imitation learning (IL)  [52, 53], where an agent learns a policy by leveraging expert demonstrations [7, 36, 37].
@@
However, in many cases, we do not have access to a live expert to query, but do have access to an abundance of logged data collected from experts. One approach to make use of this data is through offline reinforcement learning. Offline RL [31, 48] learns a policy solely from such a fixed dataset of pre-collected experiences, without the need to directly interact with the environment. Despite its advantages, offline RL often results in a suboptimal policy due to dataset limitations. This has motivated recent work that combines offline and online RL, whereby learning begins from a logged dataset before transitioning to online interactions for further improvement. While beneficial, contemporary offline-to-online RL methods suffer from catastrophic forgetting, where previously learned knowledge is overwritten during online fine-tuning, leading to significant performance degradation [40, 72].
@@
More recently, methods that integrate online RL with offline datasets utilize off-policy techniques to incorporate offline data while learning online [2, 60], mitigating catastrophic performance drops. These techniques do not require any preliminary offline RL training or incorporate specific imitation clauses that prioritize pre-existing offline data. Notably, RLPD [2] exhibits strong empirical performance, however it employs a uniform random sampling strategy for both offline and online learning, ignoring that different transitions contribute differently to the various stages of policy improvement. Furthermore, this uniform sampling strategy may lead to data inefficiencies (e.g., sampling unhelpful transitions while missing out on valuable ones) and makes policy improvement highly sensitive to data quality.
Figure 1: $A^{3}$RL combines online and offline RL using a priority-based sampling strategy that prioritizes samples from online roll-outs and offline data that align with directions for policy improvement.
@@
Our contributions. In this work, we introduce Active Advantage-Aligned Reinforcement Learning ($A^{3}$RL), a novel method that operates in the realm of online RL with an offline dataset, as illustrated in Fig. 1. Our approach dynamically prioritizes transitions with the highest potential to drive policy improvement, while also tackling the mixed offline-online nature by aligning the sampling distribution to the policy’s data generation. More specifically, $A^{3}$RL considers not only the relevance of the data in facilitating the current policy’s online exploration and exploitation, but also its estimated contribution to policy improvement via confidence-aware advantage-based prioritization. $A^{3}$RL demonstrates robustness to data quality in a black-box manner and maintains resilience under varying environmental conditions. Notably, it also effectively accelerates policy improvement, even in a purely online environment.
In short, our contributions are:
@@
  • We propose $A^{3}$RL, a novel algorithm for online RL with offline data. This algorithm surpasses current state-of-the-art (SOTA) methods through a priority-based sampling strategy integrating a conservative estimate of the advantage function with the extent of online coverage of the offline dataset.
@@
  • In contrast to RLPD and other related works [30, 56], which lack theoretical support, this study provides theoretical insights into our active advantage-aligned sampling strategy, demonstrating superiority and its minimum improvement gap over random sampling.
@@
  • Through extensive empirical evaluations on the D4RL benchmark [12], we demonstrate that $A^{3}$RL achieves consistent and significant improvements over prior SOTA models, especially on harder Adroit tasks.
@@
  • Given the black-box nature of offline datasets, we conduct comprehensive ablation studies across a range of dataset qualities and environmental settings, including the purely online scenario, to evaluate the robustness of $A^{3}$RL. These studies consistently confirm its stable performance across diverse conditions, regardless of environmental factors or data quality.
2 Related Work
@@
Online RL with offline datasets. Several methods exist that incorporate offline datasets in online RL to enhance sample efficiency. Many rely on high-quality expert demonstrations [20, 24, 43, 50, 65, 73]. Nair et al. [44] introduced the Advantage Weighted Actor Critic (AWAC), which utilizes regulated policy updates to maintain the policy’s proximity to the observed data during both offline and online phases. On the other hand, Lee et al. [30] proposed an initially pessimistic approach to avoid over-optimism and bootstrap errors in the early online phase, gradually reducing the level of pessimism as more online data become available. Most relevant to our work is RLPD [2], which adopts a sample-efficient off-policy approach to learning that does not require pre-training. Unlike RLPD, which utilizes symmetric sampling to randomly draw from both online and offline datasets for policy improvement, $A^{3}$RL adopts a Prioritized Experience Replay [56] (PER)-style method, whereby it selectively uses data from both datasets to enhance policy performance.
@@
Prioritized experience replay. Experience replay [32] enhances data efficiency in online RL by reusing past experiences. PER  [56] introduces prioritization based on temporal difference (TD) error to ensure that miscalibrated $Q$-values are corrected more frequently, and has proven effective in a variety of settings [17, 22, 43, 46, 55, 62, 66, 70]. Alternative prioritization strategies have been explored, such as prioritizing transitions based on expected returns [21] or adjusting sample importance based on recency [11]. Eysenbach et al. [9] apply a density ratio to the reward instead of weighting the samples. These works predominantly focus on either purely online or purely offline applications of PER. In contrast, our approach uniquely integrates the strengths of both online and offline data. The most relevant studies to ours include Sinha et al. [59] that uses the density ratio between off-policy and near-on-policy state-action distributions as an importance weight for policy evaluation, and Lee et al. [30] that employs density ratios to select relevant, on-policy–like samples from offline datasets. However, existing methods fail to account for a critical factor—the potential of data (e.g., its advantageousness and informativeness) that fundamentally shapes policy improvement.
@@
Active learning in RL. Active learning has been explored in RL for data-efficient exploration [8, 10, 28, 34, 36, 37, 39]. Unlike previous approaches that focus on oracle selection [36, 37], state exploration [8, 37] or reward estimation [39], $A^{3}$RL introduces an active sampling mechanism tailored to online RL with offline data, prioritizing transitions that maximize policy improvement. We defer more details of related work to the Appendix C.
3 Preliminaries and Problem Statement
@@
We consider a discounted Markov decision process (MDP) environment [3] characterized by a tuple $M=(\mathcal{S},\mathcal{A},\mathcal{P},R,\gamma,d_{0})$, where $\mathcal{S}$ represents a potentially infinite state space, $\mathcal{A}$ is the action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is the unknown transition kernel, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in(0,1)$ is the discount factor and $d_{0}\left({s}\right)$ is the initial state distribution. The learner’s objective is to solve for the policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ that maximizes the expected sum of discounted future rewards $\mathbb{E}_{\pi}\!\left[\sum_{t=1}^{\infty}\gamma^{t}r\left({s_{t},a_{t}}\right)\right]$, where the expectation is taken over the trajectory sampled from $\pi$.
@@
Maximum entropy RL. In this work, we adopt off-policy soft actor-critic (SAC) [15] RL to train an agent with samples generated by any behavior policy. We use a general maximum entropy objective [2, 15, 74] as follows:
$\max_{\pi}\mathbb{E}_{s\sim{\mathcal{R}},{a\sim\pi}}\!\left[\sum_{t=0}^{\infty}\gamma^{t}\left({r_{t}+\alpha\mathcal{H}\left({\pi\left({a|s}\right)}\right)}\right)\right],$ (1)
@@
where $\mathcal{R}$ is replay buffer and $\alpha$ is a temperature parameter. This involves optimizing reward while encouraging exploration, making the learned policy more robust.
@@
$Q$-value and advantage function. The $Q$-value function measures the expected return of executing action $a$ in state $s$ under policy $\pi$, and satisfies the Bellman self-consistency condition: $Q^{\pi}\left({s,a}\right)=\mathcal{B}^{\pi}Q^{\pi}\left({s,a}\right)$, where $\mathcal{B}^{\pi}$ is the Bellman operator: $\mathcal{B}^{\pi}Q\left({s,a}\right):=r\left({s,a}\right)+\gamma\mathbb{E}_{s^{\prime}\sim P\left({\cdot|s,a}\right)}\!\left[V^{\pi}\left({s^{\prime}}\right)\right]$. The soft state value function is defined as: $V^{\pi}\left({s}\right):=\mathbb{E}_{a\sim\pi\left({\cdot|s}\right)}\!\left[Q^{\pi}\left({s,a}\right)-{\log\pi\left({a|s}\right)}\right]$. For a policy $\pi$, the advantage function [61] quantifies the relative benefit of selecting $a$ over the policy’s default behavior:
@@
Specifically, SAC learns a soft Q-Function, denoted as $Q_{\theta}\left({s,a}\right)$ with parameters $\theta$, and a stochastic policy $\pi_{\phi}$ parameterized by $\phi$. The SAC method involves alternating between updates for the critic and the actor by minimizing their respective objectives [30] as follows
@@
$-\gamma{\mathbb{E}_{a^{\prime}_{t+1}\sim\pi_{\phi}}\!\left[Q_{\theta^{\prime}}\left({s_{t+1},a^{\prime}_{t+1}}\right)-\alpha\log\pi_{\phi}\left({a^{\prime}_{t+1}|s_{t+1}}\right)\right]})^{2}],$
@@
Here, $\mathcal{R}$ is an experience replay buffer of either on-policy experience [61] or through off-policy experience [42, 47], and $\overline{\theta}$ denotes the delayed target parameters.
@@
Prioritized experience replay. PER [56] serves as the basis of our sampling techniques, providing a framework for prioritizing experience replay based on transition importance. Instead of sampling uniformly from the replay buffer $\mathcal{R}$, PER assigns higher probability to transitions with miscalibrated $Q$-values, leading to improved sample efficiency [17] as these transitions are seen and corrected more frequently. Each transition $\mathcal{R}_{i}=(s_{i},a_{i},r_{i},s_{i+1})$ is assigned a priority $P_{i}$, typically based on the absolute of the TD-error $\delta=r+\gamma Q_{{\theta^{\prime}}}(s_{t+1},a^{\prime}_{t+1})-Q_{\theta}(s_{t},a_{t})$ where $a^{\prime}_{t+1}\sim\pi(\cdot|s_{t+1})$ [4, 17, 46, 56, 64]. Subsequently, the sampling approach of PER involves sampling an index set $\mathcal{I}$ within the range $\left[|\mathcal{R}|\right]$ based on the probabilities $p_{i}$ assigned by the priority set as follows: $p_{i}=\frac{P_{i}^{\alpha}}{\sum_{k\in\left[|\mathcal{R}|\right]}P_{k}^{\alpha}}$, with a hyperparameter $\alpha>0$. To correct for sampling bias, PER applies importance sampling weights:
$u_{i}\propto\bigl({1}/(|\mathcal{R}|\cdot p_{i})\bigr)^{\beta_{priority}},$ (3)
@@
where $\beta_{priority}$ anneals from $\beta_{0}\in(0,1)$ to $1$ during training to counteract bias in the learning updates, and the importance sampling weights are normalized to have maximum weight of $1$ for stability. While standard PER prioritizes TD-error, our method extends this framework to prioritize transitions based on onlineness and estimated advantage.
@@
Online RL with offline datasets. In this work, we study online RL, with offline datasets denoted as $\mathcal{D}$ [2]. These datasets consist of a set of tuples $(s,a,r,s^{\prime})$ generated from a specific MDP. A key characteristic of offline datasets is that they typically offer only partial coverage of state-action pairs. In other words, the set of states and actions in the dataset, denoted as $\left\{\left({s,a}\right)\in\mathcal{D}\right\}$, represents a limited subset of the entire state space and action space, $\mathcal{S}\times\mathcal{A}$. Moreover, learning on the data with incomplete coverage of state-action pairs potentially results in excessive value extrapolation during the learning process for methods using function approximation [14]. Our model, based on SAC [15], incorporates several effective strategies for RL with offline data, as outlined in RLPD [2], with the most crucial one being Clipped Double Q-Learning. The maximization objective of Q-learning and the estimation uncertainty from value-based function approximation often leads to value overestimation [63]. To address this problem, Fujimoto et al. [13] introduced Clipped Double Q-Learning (CDQ) as a means of mitigation. CDQ involves taking the minimum from an ensemble of two Q-functions for computing TD-backups. The targets for updating the critics are given by the equation $y=r\left({s,a}\right)+\gamma\min_{i=1,2}Q_{\theta^{\prime}_{i}}\left({s^{\prime},a^{\prime}}\right)$, where $a^{\prime}\sim\pi\left({\cdot|s^{\prime}}\right)$. See Appendix B for more strategies to stabilize training.
4 Algorithm
Here, we describe our active advantage-aligned sampling algorithm that is theoretically grounded in the performance difference lemma §5. At a high level, we do active learning in that we select relevant offline transitions that align better with the data distribution generated by the current policy through a density term, while also incorporating the advantage, which measures the potential impact of the transition on policy improvement.
Active density term.